### v1

# Evaluating the Fitness of Ontologies for the Task of Question Generation

Abstract
Ontology-based question generation is an important application of semantic-aware systems that enables the creation of large question banks for diverse learning environments. The effectiveness of these systems, both in terms of the calibre and cognitive difficulty of the resulting questions, depends heavily on the quality and modelling approach of the underlying ontologies, making it crucial to assess their fitness for this task. To date, there has been no comprehensive investigation into the specific ontology aspects or characteristics that affect the question generation process. Therefore, this paper proposes a set of requirements and task-specific metrics for evaluating the fitness of ontologies for question generation tasks in pedagogical settings. Using the ROMEO methodology, a structured framework for deriving task-specific metrics, an expert-based approach is employed to assess the performance of various ontologies in Automatic Question Generation (AQG) tasks, which is then evaluated over a set of ontologies. Our results demonstrate that ontology characteristics significantly impact the effectiveness of question generation, with different ontologies exhibiting varying performance levels. This highlights the importance of assessing ontology quality with respect to AQG tasks.
Keywords:
Ontology Evaluation Question Generation Ontologies
1 Introduction
Research on the use of semantic models within Automatic Question Generation (AQG) has greatly advanced the field of educational assessment. Ontologies, with their structured representation of knowledge, introduce an additional layer of semantic understanding which allows AQG systems to generate questions modelled at a semantic level [7], thereby surpassing syntax- and vocabulary-focused approaches that operate at a shallower level [17]. Furthermore, semantic-aware AQG models that exploit ontological elements such as classes and properties, have demonstrated superior generalisability across different domains and question formats, compared to machine learning-based approaches [3]. Factors such as the size of the ontology, its hierarchical organisation, domain coverage, and defined semantic relations play a key role in determining the properties of the output [23, 37, 44]; e.g., the linguistic issues in the questions generated have been found to be linked to how concepts are modelled within the ontology [23]. Yet there has been no systematic investigation into the type of ontology features (or how they are quantified) that can enhance or hinder pedagogical question generation.
The choice of ontology is crucial when generating and evaluating questions using AQG systems. Although it is common practice within the AQG research community to develop experimental ontologies [2, 5, 13, 24, 40, 42, 43], this has been criticised for potentially affecting the generalisability and validity of the frameworks and their question patterns, due to them being purpose-built and not representative of real-world applications [5]. Reusing existing ontologies would address these concerns by ensuring the frameworks had been tested on widely accepted and standardised knowledge structures, thereby enhancing their applicability and reliability across diverse contexts. However, identifying and selecting suitable ontologies that are fit for purpose is complex, as this requires the expertise, intuition, and domain knowledge of an ontology specialist. Similarly, when creating new ontologies, frameworks are needed that can guide construction and be used to evaluate the final product. Approaches such as Application or Task-based evaluation can assess the suitability and fitness of an ontology for a specific task, by evaluating the extent to which an ontology conforms to the task requirements [9, 18, 19, 26, 33]. However, for this to be applied to AQG, the requirements (and evaluation metrics) for question generation must first be specified, before they can be used to evaluate how well an ontology supports them.
This paper focuses on identifying the requirements and ontology evaluation metrics specifically designed to assess the suitability of an ontology for AQG. By using the ROMEO methodology [46] (a requirements-oriented framework designed for evaluating ontologies in task-specific contexts), a core set of requirements and associated metrics have been identified and evaluated over a set of ontologies. This expert-led approach combines theoretical and practical perspectives, by integrating domain ontologies into a question-generation model and deriving relevant metrics based on observed performance. Section 2 provides an overview of AQG together with the challenges involved in evaluating ontologies, and the methodology adopted to determine a set of targetted evaluation requirements for AQG is presented in Section 3. The resulting requirements, evaluation metrics and evaluations are detailed in Section 4, before concluding in Section 5.
2 Background and Related Work
Ontology-based question generation, whereby a domain ontology is used as the knowledge source to generate questions, typically involves the modelling of generic question types using a set of templates containing specific Resource Description Framework (RDF) patterns. These are used to generate different types of questions using ontological elements such as concepts [6, 13], various types of properties [37] and individuals [13]. Several of the most common question formats are illustrated in Table 1, together with the corresponding RDF pattern requirements for each question type (which vary both in terms of the number of triple patterns they require and the types of elements they include). Papasalouros & Chatzigiannakou [29] classify ontology-based questions based on the type of ontological components targetted by the generation strategies: class-based, property-based and terminology-based (Table 1). Class membership questions assess the learner’s understanding of class-instance relationships, such as identifying the class to which a specific entity belongs [1, 5, 13], whereas Property-based questions focus on testing the learner’s knowledge of an entity’s properties and their values, such as the attributes or relationships of a given entity [1, 30, 36]. Terminology-based questions explore hierarchical relationships of an entity within the ontology, such as subclass and superclass associations [5, 11, 29]. In addition to these question types, Annotation-based questions [5, 36] leverage annotations in the ontology to generate questions about descriptions or the definitions of terms. These generation strategies result in questions that have a one-to-one mapping between the number of relevant triples and questions generated, and as such each pattern template only tests the learners’ knowledge about a single fact.
Table 1: Ontology-based question generation patterns and approaches. Uppercase characters represent classes, whereas lowercase ones are instances. $P$ refers to properties.
Question Generation Strategy RDF Pattern Approaches
Class membership <x> <rdf:type> <X> [1, 5, 13]
Property-based <x> <P> <y> [1, 11, 30, 36]
Terminology-based <x> <rdfs:subClassOf> <y> [5, 11, 29]
Annotation-based <X> <rdfs:comment> <string> [5, 36]
MCQs <x> <rdf:type> <X> [1, 5, 13, 40]
<y> <rdf:type> <Y>
<X>, <Y> <rdfs:subClassOf> <Z>
Multi-entity <x> <P1> <y> [25, 35, 37]
<x> <P2> <z>
Beyond these foundational types, other common question types incorporate additional triples or target special types of semantic relationships, such as Multiple Choice Questions (MCQs) and multi-entity questions. MCQs require not only the generation of a question but also the creation of a set of distractors that are plausible yet distinct from the correct answer [2, 5, 13, 25, 37, 41]. Multi-entity questions synthesise knowledge from multiple interconnected entities, requiring the learner to have a deeper understanding of the relationships and dependencies between various ontological elements [25, 35, 37]. These advanced question types are particularly valuable for evaluating higher-order thinking and comprehension [12, 25, 37], and whilst MCQs and multi-entity questions may overlap with the previous categories in their use of classes, properties, or instances, their defining characteristic is the integration of multiple axioms to form complex questions and the construction of distractors, that introduces an additional cognitive challenge.
Domain ontologies, which represent the knowledge of a specific subject area, form the primary input for question generation models. Many studies rely on hand-crafted ontologies developed specifically for evaluation purposes [2, 5, 13, 24, 40, 42, 43], which frequently focus on the knowledge of a particular course [5, 24], or represent broader knowledge of a general domain [2, 13, 24, 40, 42, 43]. Consequently, such ontologies tend to be small in scope and shallow in structure, capturing only the most general aspects of the domain. This limited depth and scale may constrain the capabilities of question generation models and hinder a comprehensive assessment of their effectiveness. Additionally, the process of building domain-specific ontologies from scratch is time-intensive and requires specialised expertise, posing a challenge to the acceleration of research in this area. Furthermore, whether a new ontology is engineered or an existing one is reused, it is still important to quantitatively assess if it is fit for the task of question generation through the use of appropriate ontology evaluation metrics.
Ontology evaluation is a long-standing research area that has evolved together with the growth of ontology engineering and semantic web technologies. Early studies focused on formal evaluation methods that consider aspects such as logical consistency [16], and more recently, broader and more practical considerations are considered such as domain relevance, user-centric adaptability and performance within specific applications [33, 34].
Ontologies, as complex knowledge artefacts, can be evaluated across multiple layers to manage their complexity [9]:
  • lexical/vocabulary: examines the modelled concepts and their representation;
  • hierarchy/taxonomy: focuses on subsumption relationships;
  • semantic: checks non-taxonomic relations;
  • application: assesses the ontology’s performance and fitness for specific tasks;
  • syntactic: verifies its adherence to the representation language’s rules;
  • design: ensures compliance with predefined design principles.
As ontologies are used across diverse application contexts, it is crucial that robust evaluation methods are used to assess them before their adoption within new systems. Ontology evaluation involves a comprehensive examination of their structure, alignment with domain requirements, and use for specific tasks [18, 19]. Such methods can be broadly categorised into four main categories [9, 19, 45]:
Gold Standard evaluation:
where the ontology is compared against a gold standard, which is an established and authoritative ontology or knowledge base covering the same domain. This evaluates the ontology’s coverage by identifying gaps, inconsistencies, or redundancies.
User-centred evaluation:
this involves obtaining and analysing feedback from end-users or domain experts who evaluate the ontology.
Data-driven evaluation:
directly compares an ontology to existing data, such as textual documents related to the domain it models.
Application or task-based evaluation:
this is performed when an ontology is used for a specific application or task. This evaluation assesses the effectiveness of the ontology in the context of an application or task.
When reusing an ontology for a specific task, it is imperative to evaluate its suitability and alignment with the task’s requirements. A requirement is defined as “an expression of desired behaviour[32], with ontology requirements specifically denoting the desired quality or competency of an ontology within the context of an application [46]. Given the diversity of ontology requirements across applications, a thorough analysis of the target application needs is essential before ontology reuse. Once these requirements are identified, they can be operationalised qualitatively, through expert judgments or user feedback, or quantitatively, employing measurable metrics (i.e. measures). Although various evaluation measures for ontologies have been proposed, selecting appropriate measures should be guided by the application’s requirements rather than relying on an arbitrary set of metrics [16, 27, 31, 46].
3 Requirement-Oriented Methodology for AQG
In order to determine an appropriate set of ontology-evaluation criteria for AQG, it is necessary to first identify those ontology characteristics that are both relevant to the generation of good quality questions, as well as being measurable. ROMEO (Requirements-Oriented Methodology for Evaluating Ontologies) is a requirement-oriented approach for identifying a set of unique requirements and characteristics for a task domain (in this context, AQG), that can then be aligned to a set of relevant evaluation metrics [46]. It was originally adapted from the Goal-Question-Metrics (GQM) framework [10], designed to derive quality measures by linking them to the goals of a given software application. By determining the requirements of the AQG process through a user-based study, suitable metrics can be identified that quantify the degree to which an ontology can effectively fulfil its intended purpose; i.e. facilitating the generation of high calibre and cognitively challenging questions.
An expert-led study was conducted to determine a number of ontology requirements given a set of generated questions, based on the following stages:
  1. 1.
    A small number of existing domain ontologies were selected for the purpose of generating questions (Table 2); namely, an astronomy-based ontology focused on the solar system,111 https://anonymous.4open.science/r/SolarSystemOntology-F64E/ an ontology representing U.S. geographical data, [^2] http://www.cs.utexas.edu/users/ml/geo.html and an ontology capturing intuitive knowledge about African wildlife [20].
  2. 2.
    Several question generation patterns were utilised to generate questions by instantiating templates with ontology elements (Table 1). This resulted in the generation of 3482 questions [^3] For MCQs, only unique question stems were considered; repeated questions with varying answer choices were excluded to avoid redundancy. across all three ontologies (Table 4).
  3. 3.
    An expert familiar with AQG was asked to assess the quality of questions, by completing a questionnaire containing a mix of open and closed questions. Two criteria were used to assess the questions; effectiveness and appropriateness (see below). Furthermore, they were asked to identify any ontology-related requirements that may impact the quality of the generation process.
  4. 4.
    The expert responses were qualitatively analysed to gain a comprehensive understanding of the expert’s insights, resulting in a set of requirements (discussed in Section 4) that were then used to define the relevant metrics.
The ontologies listed in Table 2 were chosen as they exhibit different characteristics and thus could be used to examine how their structural differences influence the question generation outcome; such as an emphasis on taxonomic modelling or hierarchical classifications and conceptual relationships. Two of these were also populated, thus providing rich factual and instance-based representations. The questions that were presented to the expert were generated using an AQG approach [4] that utilised the RDF patterns in Table 1 within a fixed textual template. Additional post-processing techniques, such as verbalisation (which are typically carried out to enhance the generation results), were intentionally excluded to maintain consistency in the structure and format of the generated questions across different ontologies, and to ensure that the expert-led comparison of question quality and effectiveness was not influenced by such techniques. [^4] Example questions appearing later in this paper have been paraphrased for clarity.
Table 2: Basic statistics of the ontologies utilised in study.
Ontology Axioms Classes Number of Properties Individuals
Object Datatype Annotation
Solar System 328 10 3 7 1 70
Geography 3573 9 17 11 3 713
African Wildlife 108 31 5 0 1 0
An expert reviewed the question specifications detailed in Table 1 and evaluated the questions generated from each of the three ontologies; with the aim of identifying the specific ontological features that contributed to the variations in the questions. The two evaluation criteria employed by Leo et al. [25] were used when assessing each ontology and its corresponding questions:
  • The effectiveness of the questions is assessed by considering: 1) the number and variety of questions generated; 2) addressing whether the ontology produced all question types; and 3) comparing the quantity of questions across ontologies and templates. This is done by analysing the questions in relation to defined ontological relationships.
  • The question quality (appropriateness) is assessed based on how pedagogically appropriate the questions are in order to assess learners at different knowledge levels, by ensuring that there is a range of questions from basic to more advanced understanding of the domain.
As the evaluation task involved an extensive qualitative analysis of a large pool of questions generated from three diverse ontologies, a single expert was recruited by this study to assess the feasibility of the proposed approach, as well as to understand critical aspects such as the time and effort needed by experts for such evaluation.
4 Response Analysis and Requirement Derivation
This section presents the outcomes of the expert-led study, by describing the observations made by the expert, the requirements derived from these observations, and proposed metrics that measure the requirements. The proposed evaluation metrics build upon well-established measures [16, 28, 38, 39] that have been adapted to align with the specific requirements of the AQG task. The model used for defining metrics formally specifies the ontology schema and the knowledge base structures. An ontology schema represents the class and property definitions (i.e. T-box) and is denoted by the 5-tuple $\mathcal{O}=\{\mathcal{C},\mathcal{P},\mathcal{H}^{\mathcal{C}},prop,\mathcal{A}\}$, whereas a knowledge base (i.e. A-box) represents the instantiated ontological elements (individuals) and is denoted by the 5-tuple $\mathcal{KB}=\{\mathcal{O},\mathcal{I},L,inst^{\mathcal{C}},inst^{\mathcal{P}}\}$. The meaning of the symbols used in these definitions are given in Table 3. Whilst this model is similar to others proposed for various ontology scenarios [28], it has been adapted to be more pertinent to the metrics that focus on AQG tasks.
Table 3: Symbols used by the metrics following the nomenclature used in [38, 39].
Ontology Schema
$\mathcal{O}=\{\mathcal{C},\mathcal{P},\mathcal{H}^{\mathcal{C}},prop,\mathcal{A}\}$
$\mathcal{C}$ a set of concepts, disjoint from $\mathcal{P}$. $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ denotes the subset of populated concepts (i.e., concepts that have at least one instance $\mathcal{I}$ in $\mathcal{KB}$).
$\mathcal{P}$ a set of properties, disjoint from $\mathcal{C}$.
$\mathcal{H}^{\mathcal{C}}$ a concept hierarchy, which is a directed, transitive relation, and relates to the taxonomy. $\mathcal{H}^{\mathcal{C}}(\mathcal{C}_{1},\mathcal{C}_{2})$ states that $\mathcal{C}_{1}$ is a sub-concept of $\mathcal{C}_{2}$ (i.e. $\mathcal{C}_{1}\sqsubseteq\mathcal{C}_{2}$).
$prop$: $\mathcal{P}\to\mathcal{C}\times\mathcal{C}$: A function mapping properties that link pairs of concepts. $prop(\mathcal{P})=(\mathcal{C}_{1},\mathcal{C}_{2})$ can also be expressed as $\mathcal{P}(\mathcal{C}_{1},\mathcal{C}_{2})$
$\mathcal{A}$ a set containing annotation properties of the type rdfs:comment
Knowledge Base
$\mathcal{KB}=\{\mathcal{O},\mathcal{I},L,inst^{\mathcal{C}},inst^{\mathcal{P}}\}$
$\mathcal{O}$ the Ontology Schema.
$\mathcal{I}$ a set of instance identifiers disjoint from $\mathcal{C}$ and $\mathcal{P}$.
$L$ a set of literal values.
$inst^{\mathcal{C}}$: $\mathcal{C}\to 2^{\mathcal{I}}$ (concept instantiation) maps concepts to their instances. $inst^{\mathcal{C}}(\mathcal{C})=\mathcal{I}$ may also be written as $\mathcal{C}(\mathcal{I})$.
$inst^{\mathcal{P}}$: $\mathcal{P}\to 2^{\mathcal{I}\times Ob}$ maps properties to their instantiated objects (i.e. $Ob=\mathcal{I}\cup L$). $inst^{\mathcal{P}}(\mathcal{I},Ob)=\mathcal{I}$ may also be written as $\mathcal{P}(\mathcal{I},Ob)$.
4.1 RQ1: Include essential RDF pattern pre-requisites.
Observation: The first analysis examined whether the ontologies could generate all question types listed in Table 1. The Geography ontology was the largest considered (in terms of axioms), and consequently generated the highest number of questions (Table 4), although the majority of these were property-based. However, size alone does not determine its suitability for the task; whilst obviously influential, structural complexity is also equally (if not more critically) important in accommodating the diverse axiom requirements of different question types. For example, no annotation-based questions were generated for Geography as it lacked annotation properties that could match the pattern <x> <rdfs:comment> <string>. The Solar System ontology generated questions for each of the types, reflecting a comprehensive structure and completeness in satisfying the question pattern pre-requisites from Table 1. This contrasts with African Wildlife, which suffered from a lack of individuals, and only generated terminology-based and annotation-based questions. These observations emphasise that structural complexity, including the presence of all prerequisites (e.g. the patterns in Table 1) are essential for generating a comprehensive range of question types. Missing axioms or incomplete structures can significantly limit the ontology’s ability to support specific templates, hindering its effectiveness for question generation.
Requirement: An input ontology should assist in generating all desirable question types, regardless of the axiom prerequisites required for each question. This requirement can be evaluated through metrics that address two key questions:
  1. 1.
    “How many RDF pattern fragments align with the components of the ontology schema and metadata model?”
  2. 2.
    “What is the structural complexity level of each component in the ontology?”
Table 4: Number of questions generated for each question generation strategy.
Solar System Geography African Wildlife
Class membership 70 713 0
Property-based 120 2044 0
Terminology-based 9 1 25
Annotation-based 10 0 15
MCQs 8 0 0
Multi-entity 25 724 0
Total 242 3482 40
Metric Pattern Coverage ($PC\in[0,1]$): The first question assesses the completeness of the ontology based on its utilisation of RDF pattern fragments (Table 1), described by the instantiation of concepts ($inst^{\mathcal{C}}$) and objects ($inst^{\mathcal{P}}$), the class hierarchy ($\mathcal{H}^{\mathcal{C}}$) and schema annotations ($\mathcal{A}$). The metric, defined as $PC=\mathcal{F}_{\text{used}}/\mathcal{F}_{\text{tot}}$ quantifies the ratio of pattern fragments used ($\mathcal{F}_{\text{used}}$) to the total available defined within the AQG patterns ($\mathcal{F}_{\text{tot}}$), indicating how effectively the ontology leverages its formal language. A high score suggests full pattern utilisation, while a lower score identifies missing elements that may hinder question generation (see Table 5, together with the scores for each ontology evaluated).
Solar System achieved full coverage ($PC=1$), aligning with its ability to generate all question types (Table 4). Geography scored lower due to missing pattern types, particularly the absence of annotation properties (rdfs:comment), leading to $\mathcal{A}=\emptyset$. African Wildlife covered only half of the patterns, limiting its question generation to taxonomic and annotation types.
Table 5: Comparative evaluation of ontologies across the proposed evaluation metrics.
Evaluation Name Metric Solar Geography African
System Wildlife
Pattern Coverage $PC=\tfrac{\mathcal{F}_{\text{used}}}{\mathcal{F}_{\text{tot}}}$ 1 0.7 0.5
Class Richness $CR=\tfrac{|\mathcal{C}^{\prime}|}{|\mathcal{C}|}$ 0.8 1 0
Average Population $P=\tfrac{|\mathcal{I}|}{|\mathcal{C}|}$ 7 79 0
Inheritance Richness $IR=\tfrac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}\left|\mathcal{H}^{\mathcal{C}}(\mathcal{C}_{1},\mathcal{C}_{i})\right|$ 0.9 0.1 1.2
Relationship Diversity $RD=\tfrac{|inst^{\mathcal{P}}|}{|\mathcal{H}^{\mathcal{C}}|+|inst^{\mathcal{P}}|}$ 0.9 0.9 0
Relationship Richness $\overline{RR}=\tfrac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}RR_{C}$ 0.8 0.9 0
Average Connectivity $\overline{Cn}=\tfrac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}Cn(\mathcal{C}_{i})$ 4.8 216 0
Sibling fan-outness $SF=\frac{|C^{\prime}_{\text{sib}}|}{|C|}$ 8 0 0
Average Depth $\overline{D}=\frac{1}{|\mathcal{C}|}\sum_{\mathcal{C}_{i}\in\mathcal{C}}CD(\mathcal{C}_{i})$ 2.3 1.1 2.3
Metric Class Richness ($CR\in[0,1]$): The second question (“What is the structural complexity level of each component in the ontology?”) extends the first by assessing structural complexity at both schema and data levels through two key metrics: Class Richness and Average Population (adopted from Tartir et al. [38, 39]). Class Richness ($CR=|\mathcal{C}^{\prime}|/|\mathcal{C}|$) measures the ratio of classes in the ontology ($\mathcal{C}$) that have been instantiated within the Knowledge Base ($\mathcal{C}^{\prime}$); which is essential for generating class-based questions (Table 5). A low $CR$ indicates a conceptually rich schema but insufficient KB instances for question generation.
Ontologies with well-populated KBs, such as Solar System and Geography, achieved high $CR$, while African Wildlife, lacking instances, scored zero.
Metric Average Population ($P\in\mathbb{R}$): measures the average number of instances per class, indicating whether classes are sufficiently populated ($P=|\mathcal{I}|/|\mathcal{C}|$). This directly affects class-based question generation and other instance-dependent templates, such as property-based strategies. Together with $CR$, it reveals the ontology population distribution, influencing both question diversity and volume.
For example, Geography (79 instances per class) generated 713 questions, whereas Solar System (7 instances per class) produced only 70.
Metric Inheritance Richness ($IR\in\mathbb{R}$): As terminology-based questions are solely dependent on the taxonomic relationships defined within the ontology, it is necessary to quantify the ratio between the total number of subclasses ($|\mathcal{C}_{1}|$) and the total number of classes $|\mathcal{C}|$ defined in the ontology schema. This metric determines the average number of subclasses per class ($IR={|\mathcal{C}|}^{-1}\sum_{\mathcal{C}_{i}\in\mathcal{C}}\left|\mathcal{H}^{\mathcal{C}}(\mathcal{C}_{1},\mathcal{C}_{i})\right|$).
African Wildlife achieved the highest score ($IR=1.2$) reflecting the higher number of terminology-based questions generated overall (Table 4).
Metric Relationship Diversity ($RD\in[0,1]$): To evaluate the ontology’s capability to generate property-based questions, two additional metrics are examined: Relationship Diversity and Relationship Richness. This question generation strategy relies on the use of properties $\mathcal{P}$ to connect objects $Ob=\mathcal{I}\cup L$ (i.e. instances $\mathcal{I}$ or literals $L$) thus targeting the instantiated properties $inst^{\mathcal{P}}$ in the A-Box, resulting in the formation of more detailed questions regarding concepts’ properties. Relationship Diversity reflects the balance between taxonomic and non-taxonomic relations, and is defined as $RD=|inst^{\mathcal{P}}|~{}/~{}(|\mathcal{H}^{\mathcal{C}}|+|inst^{\mathcal{P}}|)$. A high $RD$ indicates an emphasis on non-taxonomic relations, increasing property-based questions while reducing terminology-based ones.
Solar System and Geography scored highly, favouring property-driven questions. However, African Wildlife ($RD=0$) prioritised taxonomic knowledge but, as it is lacking in instances, generated no questions of either type (Table 4).
Metric Relationship Richness ($\overline{RR}\in\mathbb{R}$): This quantifies the number of properties defined in the schema (T-box) for each class, that are utilised by instances in the KB (A-box) by finding the mean relationship richness for each class ($RR_{C}$). This is defined as the ratio of the number of properties $\mathcal{P}$ used by the instances $\mathcal{I}_{i}$ of class $\mathcal{C}_{i}$ (i.e. $|\mathcal{P}(\mathcal{I}_{i},Ob)|$) and the total number of properties defined for class $\mathcal{C}_{i}$ in the schema (i.e. $|\mathcal{P}(\mathcal{C}_{i},\mathcal{C}_{j})|$):
[…]
An ontology with a high $\overline{RR}$ (e.g. $\overline{RR}=0.8$ and $\overline{RR}=0.9$ for Solar System and Geography respectively) indicates a well-structured axiomatic foundation aligned with the prerequisites, highly suited for generating property-based questions.
4.2 RQ2: Supports the generation of multi-entity questions.
Observation: The Geography ontology generated the largest number of multi-entity questions (724 questions); e.g. “What is the highest point in Georgia with an elevation of 1,458 meters?”. The expert attributed this to the rich and diverse range of descriptions within that ontology relating to the breadth of properties used to model each concept; for example, the concept State in Geography included the properties statePopulation, stateArea and borders. Likewise, Solar System, though much smaller in size, generated 25 multi-entity questions, again due to having a broad set of properties. However, no multi-entity questions were generated for African Wildlife, as it lacked instances in its KB.
Requirement: The effectiveness of an ontology should consider support for the generation of the more advanced multi-entity questions. This type of question involves the use of more than one axiom, allowing for greater variation in question generation, thus testing learners’ knowledge of how concepts interact [37]. This variation is also important as it enables the creation of questions that go beyond simple, one-fact questions, offering a broader and deeper understanding of the domain (i.e. targetting the second level in Bloom’s taxonomy [8, 37]). Multi-entity questions tend to be more detailed and complex, and cover a wider breadth of knowledge within the domain. These questions are generated when an ontology contains rich concept descriptions, with multiple axioms that comprehensively define a concept by detailing its properties and associations [37].
Metric Average Connectivity ($\overline{Cn}\in\mathbb{R}$): The fact that multi-entity questions are formed by combining multiple facts regarding a concept poses the question “How well are instances of each class connected to objects?” Thus, a metric for this requirement should relate to Connectivity ($Cn$), which quantifies the out-degree (i.e. number of properties to other concepts) of a concept. This metric captures the centrality of classes within the ontology by quantifying how connected a class is to other objects ($Ob$), and is defined for properties $\mathcal{P}$ as:
[…]
Thus, at the ontology level, Average Connectivity ($\overline{Cn}$) measures the overall degree of interconnectivity across all classes in the ontology.
Geography scores highly ($\overline{Cn}=216$) and Solar System less so but with a positive score ($\overline{Cn}=5$), suggesting that both are well-suited for generating multi-entity questions, which typically involve 2–3 connections per question.
4.3 RQ3: Supports the generation of plausible MCQ distractors.
Observation: Generating MCQs posed challenges across all three ontologies. Questions of the form “Which of these is X?” typically list the answer choices with the correct answer as a subclass of $X$ and incorrect answers as sibling classes of $X$. No MCQs were generated for African Wildlife, due to a lack of instances within the sibling classes, whereas the reason for the lack of MCQs for Geography was that no sibling classes were identified. Solar System contains a rich underlying knowledge graph with each class having several sub-classes, and thus well-defined sibling relationships. Of the three ontologies, this was the only one that successfully generated MCQs.
Requirement: Multiple Choice Questions (MCQs) are the most common question types generated as they offer a more complex structure that requires the incorporation of several axioms with special characteristics. The structure of an ontology plays a key role in this process, particularly in the generation of plausible distractors for MCQs. These distractors are crafted to resemble the correct answer by leveraging semantic relationships within the ontology. Specifically, sibling classes (i.e. those that share a common immediate parent) are ideal candidates for distractors due to their inherent semantic similarity [13, 30]. In order to generate plausible distractors, an ontology should contain a hierarchical structure where populated sibling classes are well-represented, leading to the question “Does the ontology contain populated classes with a sufficient number of sibling classes?”.
Metric Sibling Fan-Outness ($SF\in\mathbb{R}$): This can be measured by calculating the ratio of the absolute populated sibling cardinality ($\mathcal{C}^{\prime}_{\text{sib}}$) to the total number of populated classes ($\mathcal{C}^{\prime}$) in the ontology; i.e. $SF=|C^{\prime}_{\text{sib}}|/|C|$. A high $SF$ score indicates that the ontology has a horizontally branching structure with multiple sibling classes under a few superclasses. This results in a broad hierarchy where categories are general rather highly differentiated. Such a structure is particularly effective for generating MCQ distractors with close semantic similarity.
Solar System was the only ontology to exhibit this pattern and successfully generated MCQs ($SF=8$), whereas as Geography lacked classes that shared a common parent, no MCQs were generated. Likewise, as African Wildlife lacked any instances in its KB, no MCQs were generated ($SF=0$).
4.4 RQ4: Supports the generation of questions with varied conceptual diversity.
Observation: An analysis of question quality with a focus on cognitive challenge considered how questions varied with respect to the level of knowledge required to successfully answer the question as evaluated by the expert. While several factors affect cognitive complexity, the emphasis here is on ontology-driven factors rather than external complexities such as the domain itself. When examining the types of questions generated from taxonomic relationships, the expert noted that Geography generated questions about high-level domain concepts (e.g. City and River). In contrast, African Wildlife and Solar System generated questions targeting both broad and more specific concepts. Kurdi [23] demonstrated that the diversity of generated questions depends not only on the size of the ontology but also on the richness of its hierarchical structure.
Requirement: A well-structured ontology should support conceptual diversity by enabling question generation across different levels of understanding, from fundamental to complex. This ensures broad domain coverage and facilitates the generation of cognitively diverse questions. Several studies have also considered conceptual depth as an important characteristic that directly impacts the ability to create cognitively diverse questions [14, 15, 21, 22, 40, 42].
Metric Average Depth ($\overline{D}\in\mathbb{R}$): This requirement can be assessed by considering whether the ontology includes concepts from varying depths of the hierarchical structure. To evaluate this, $\overline{D}={|\mathcal{C}|}^{-1}\sum_{\mathcal{C}_{i}\in\mathcal{C}}CD(\mathcal{C}_{i})$ calculates the average depth $CD(\mathcal{C}_{i})$ of each concept $\mathcal{C}_{i}\in\mathcal{C}$, defined as the number of edges from $C_{i}$ to the root in hierarchy $H^{c}$. This metric (similar to that defined by Gangemi et al. [16]) ensures question generation spans general to specific concepts, highlighting that hierarchical richness, rather than size alone, drives cognitive diversity.
Both Solar System and African Wildlife (each score $\overline{D}$ = 2.3) exhibited rich hierarchies, generating questions at varying conceptual levels. For example, Solar System included broad concepts like “Planet” while deeper levels introduced distinctions such as “Terrestrial Planet” and “Gas Giant”.
5 Conclusion
In this paper, the goal was to identify metrics for evaluating the fitness of ontologies for the task of AQG. The ROMEO methodology, a structured approach for identifying metrics based on specific task requirements was employed as a methodological framework. An expert-based metric derivation process was used to assess how different ontologies perform in AQG tasks. It was found that the characteristics of ontologies significantly influence the effectiveness of the question generation process. Our findings underscore that different ontologies yield varying levels of performance, highlighting the critical need to assess ontology quality in AQG. A set of structural and knowledge-based metrics was identified for assessing the fitness of ontologies for AQG tasks. This foundational study opens avenues for future work which should focus on exploring additional metrics that consider the content and user-oriented aspects of ontologies.